# A Survey on Benchmarks of Multimodal Large Language Models

### v1

@@ 7.2 Audio
VisualCoTshao2024VisualCoT specific local region identification free-from GPT
@@ 7.4 Omnimodal
In the real world, people are surrounded by audio, images, videos, and text messages in their daily lives. These diverse modalities collectively enhance our ability to perceive and understand scenes. However, most benchmarks focus on a single modality, which limits their ability to adapt to real-world environments that involve multiple modalities. Therefore, some works focused on evaluating MLLMs’ capabilities in handling multiple modalities simultaneously. MusicAVQA li2022MusicAVQA was constructed to evaluate MLLMs’ ability to answer questions regarding visual objects, sounds, and their associations together. However, the visual scenes in MusicAVQA were limited to music performances, where the questions were only about instrument relationships, lacking exploration of more real-life scenarios. In response to this limitation, Yang et al. AVQA proposed AVQA, which was designed for audio-visual question answering on general videos of real-life scenarios. Furthermore, to cover a wider range of modal information, Chen et al. chen2024MCUB introduced a benchmark called the Multimodal Commonality Understanding Benchmark (MCUB), which included four modalities—image, audio, video, and point cloud. The task of MCUB was to measure the model’s ability to identify commonalities among input entities from diverse modalities, but overlooked the various other relationships among different modalities. To achieve general purpose multimodal intelligence, Ying et al.ying2024MMT-Bench present MMT-Bench, a comprehensive benchmark designed to assess MLLMs across massive multimodal tasks, which required expert knowledge and deliberate visual recognition, localization, reasoning, and planning.

### v2

@@ 7.2 Audio
Grounding-Benchzhang2023Grounding-Bench grounding and chat capabilities 1000 open GPT 9
@@
VisualCoTshao2024VisualCoT specific local region identification free-from GPT 3
@@
M3CoTchen2024M3CoT multi-domain multi-step chain-of-thought reasoning 11459 MQA N/A 5
@@
MuCRli2024MuCR multimodal causal reasoning 400 MQA, open N/A 17
SPARKyu2024SPARK multi-vision sensor perception and reasoning 6,248 Y/N, MQA N/A 10
MultiMedmo2024MultiMed multimodal and multitask medical understanding 2.56 million open domain-specific metrics
HR-Benchwang2024HR-Bench high-resolution image perception 200 MQA N/A 13
@@ 7.4 Omnimodal
In the real world, people are surrounded by audio, images, videos, and text messages in their daily lives. These diverse modalities collectively enhance our ability to perceive and understand scenes. However, most benchmarks focus on a single modality, which limits their ability to adapt to real-world environments that involve multiple modalities. Therefore, some works focused on evaluating MLLMs’ capabilities in handling multiple modalities simultaneously. MusicAVQA li2022MusicAVQA was constructed to evaluate MLLMs’ ability to answer questions regarding visual objects, sounds, and their associations together. However, the visual scenes in MusicAVQA were limited to music performances, where the questions were only about instrument relationships, lacking exploration of more real-life scenarios. In response to this limitation, Yang et al. AVQA proposed AVQA, which was designed for audio-visual question answering on general videos of real-life scenarios. Furthermore, to cover a wider range of modal information, Chen et al. chen2024MCUB introduced a benchmark called the Multimodal Commonality Understanding Benchmark (MCUB), which included four modalities—image, audio, video, and point cloud. The task of MCUB was to measure the model’s ability to identify commonalities among input entities from diverse modalities but overlooked the various other relationships among different modalities. To achieve general-purpose multimodal intelligence, Ying et al.ying2024MMT-Bench present MMT-Bench, a comprehensive benchmark designed to assess MLLMs across massive multimodal tasks, which required expert knowledge and deliberate visual recognition, localization, reasoning, and planning.